OpenStudy (anonymous):

Solve -(2x + 4) = 18. -11 -7 7 11

OpenStudy (anonymous):

(-2x-4)=18 -2x=22 then you just divide by two .

OpenStudy (anonymous):

distribute the negative -2x-4 =18 add 4 and -2x =22 divide by -2 x= -11

OpenStudy (ikram002p):

\[2x+4=-18\rightarrow 2x=-22\rightarrow x= -11\]

OpenStudy (anonymous):

first start with ditrubitive property.